OpenStudy (anonymous):

Matthew purchased a washer and dryer for $2,315 using an 18-month deferred payment plan with an interest rate of 24.92%. What is the balance after the deferment period if payments of $95 were made each month?

OpenStudy (anonymous):

I think it would be $512.1 2,315/24.92= 92.90 2315-92.90=222.1 95x18=1710 222.1-1710=512.1 I'm not certain that's how you do it but I think so

OpenStudy (anonymous):

Oh I figured out the answer. Thanks for helping :)